xx EOSPolytrope

EOSxPolytrope

Ian Hawke

22/4/2002

Abstract

EOSxPolytrope

1 The equations This equation provides a polytropic equation of state to thorns using the CactusEOS interface found in EOSxBase. As such it’s a fake, as EOSxBase assumes that, e.g., the pressure is a function of both density and specific internal energy. Here the pressure is just a function of the density, and is set appropriately (the specific internal energy is always ignored).

The two fluid constants are (eosxk) and (eosxgamma), which default to 100 and 2 respectively. The formulas that are applied under the appropriate EOSxBase function calls are

To calculate the units of the Cactus quantities and back, remember that in Cactus. Here is one example how to calculate densities: and one example for calculating (for ):

2 Parameters

11eosxgamma 10Scope: restrictedx01  REAL

11Description: Adiabatic Index for Ideal Fluid

11 Range 10  x01  Default: 2.0

11:  10

11

11eosxk  10Scope: restrictedx01  REAL

11Description: Polytropic constant

11 Range 10  x01  Default: 80.0

11:  10

11

11gammaxini10Scope: restrictedx01  REAL

11Description: Polytropic Gamma used for the initial model (e.g., by RNSID)

11 Range 10  x01  Default: 2.0

11:  10

11

11usexcgs 10Scope: restrictedx01  BOOLEAN

11Description: Use the CGS units

11         10  x01  Default: no

11

3 Interfaces

General Implements:

eosx2dxpolytrope

Inherits:

eosxbase

Uses header:

EOSxBase.h

EOSxBase.inc

4 Schedule This section lists all the variables which are assigned storage by thorn EinsteinEOS/EOSxPolytrope. Storage can either last for the duration of the run (Always means that if this thorn is activated storage will be assigned, Conditional means that if this thorn is activated storage will be assigned for the duration of the run if some condition is met), or can be turned on for the duration of a schedule function.

Storage NONE

Scheduled Functions

CCTKxSTARTUP x eosxpolytropexstartup x setup the polytropic eos x

11x11Language:11fortran
11 x 11Type: 11function